OpenStudy (anonymous):

help please. simplify 1√196 a. 1/14 b. 13 c. 1/12 d. 1/169

OpenStudy (jojo4eva):

a. 1/14

OpenStudy (anonymous):

can you help with a couple more? @jojo4eva

OpenStudy (deadpoolroxz):

yes

OpenStudy (jojo4eva):

all you need to do is find out what the square root of 196 is. The square root of 196 is 14 thus making it 1/14

OpenStudy (jojo4eva):

sure if you want

OpenStudy (anonymous):

evaluate x/y for x= 2/3 and y = 6/7. a. 7/9 b. 8/21 c. 8/10 d. 12/21

OpenStudy (cwrw238):

you can get the square root of 196 using your calculator

OpenStudy (cwrw238):

invetr 6/7 then multiply:-

OpenStudy (cwrw238):

* invert